# Learning Large-Scale Competitive Team Behaviors with Mean-Field Interactions and Online Opponent Modeling

### v2

@@ 4 MF-MAPPO for Partially Observable MFTGs
Similarly, we can define Lipschitz continuous policies for the Red team with constant $L_{\psi}.$ This idea of penalizing the gradient of the policies was introduced in robotics to promote smooth and stable policies in order to aid sim-to-real transfer [Chen et al.,, 2024; Shin et al.,, 2025].
@@
In lieu of Proposition 3, we propose a communication network-based decentralized estimation filter, namely, Dynamic-Projected Consensus (D-PC). It is an extension of the control-theoretic constrained consensus problem [Nedić and Liu,, 2016] and addresses the shortcomings of the estimation algorithm proposed in Benjamin and Abate, 2025b , namely, estimation under limited communication rounds, and ensuring valid estimates in the presence of errors. Following the connected graph topology used in MFGs [Benjamin and Abate, 2025b, ], we define a state-based visibility graph $\mathcal{G}_{t}^{\mathrm{viz}}$ and team-based communication graphs $\mathcal{G}_{{\rm Blue},t}^{\mathrm{com}}$ and $\mathcal{G}_{{\rm Red},t}^{\mathrm{com}}$. We also define the projection operator $\Omega_{\mathcal{R}(x)}[\eta]\triangleq\arg\min_{\omega\in\mathcal{R}(x)}\|\eta-\omega\|_{2}$ for $\eta,\omega\in\mathbb{R}^{|\mathcal{Y}|}$ where $\mathcal{R}(x)$ is a closed and convex set. We assume that all agents in the same state receive the same information, so naturally they have the same estimate, i.e., all Blue agents at state $x\in\mathcal{X}$ at time $t$ have the same estimate $\hat{\nu}^{N_{2}}_{x,t}$ of the Red team MF. We consider two time scales: $t$ for system dynamics (2) and $\tau$ for communication rounds. At time $t$, and $\tau=0$, each Blue agent at state $x\in\mathcal{X}$ holds a belief $\hat{\nu}^{\tau=0}_{x,t}$ consistent with $\mathcal{G}_{t}^{\mathrm{viz}}$. At communication round $\tau$, agents share estimates $\hat{\nu}^{\tau-1}_{x,t}$ with neighbors defined by $\mathcal{G}_{{\rm Blue},t}^{\mathrm{com}}$ and perform for $R_{\textrm{com}}$ communication rounds: a) weighted-average consensus; and b) a projection onto a closed and convex constraint set $\mathcal{R}(x)$. The set $\mathcal{R}(x)$ combines the Red team’s MF components known with certainty by the Blue agents at state xx—i.e., the observable states given by $\mathcal{G}_{t}^{\mathrm{viz}}$—with those that must be estimated. $\mathcal{R}(x)$ guarantees that operations such as information aggregation, averaging, or distributed communication do not alter the parts of the distribution that are known with certainty. See Appendix A for detailed definitions.
@@ 5 Numerical Experiments
In this section, we evaluate MF-MAPPO across large-population scenarios using our custom-made benchmark simulation platform, MFEnv. Built as an extension of Gymnasium [Towers et al.,, 2024], MFEnv is developed specifically to facilitate research in MFTGs, supporting both finite-agent simulations and oracle-based infinite-population models. Unlike existing toolkits, MFEnv includes aggregate reward metrics, policy-versus-policy evaluation, and flexible APIs for custom mean-field environments that adhere to MF dynamics, rewards and information structures. We showcase MF-MAPPO’s efficacy on two representative environments (1) a constrained-action variant of the classical rock–paper–scissors game [Raghavan,, 1994], enabling validation against analytically computed equilibria and (2) a complex battlefield setting where Blue and Red teams engage in attack–defense tasks with higher-dimensional state and action spaces, requiring sophisticated team-level coordination. Additional results and environments (not limited to ZS-MFTGs) in Appendix E.
Constrained Rock-Paper-Scissors (cRPS). The state space of each individual agent is $\mathcal{S}=\{\texttt{R,P,S}\}$, representing rock, paper, and scissors, respectively. We consider a non-trivial restriction of the action space to $\mathcal{A}=\{\texttt{CW},\texttt{Stay}\}$ allowing agents to either move clockwise ( $\texttt{R}\rightarrow\texttt{P}$, $\texttt{P}\rightarrow\texttt{S}$, $\texttt{S}\rightarrow\texttt{R}$) or remain idle, respectively. We assume deterministic transitions, where each action leads to a unique next state deterministically. At each time step $t$, the Blue (Red) team receives a team reward $r_{t}(\mu_{t},\nu_{t})={\mu}_{t}^{\mbox{\sf T}}A{\nu}_{t}$ (${\mu}_{t}^{\mbox{\sf T}}A{\nu}_{t}$) where $A$ is the standard RPS payoff matrix.
@@
Figure 2(a) compares trajectories from MF-MAPPO and DDPG-MFTG [Shao et al., 2024a, ]. Existing algorithms such as MADDPG and MAAC [Lowe et al.,, 2017] are unsuitable as baselines because their centralized action-value functions require the actions of all agents and the full state information, leading to poor scalability in the large-population settings considered in this work. We see that MF-MAPPO successfully reaches the equilibrium distribution. The transient time is attributed to finite-population approximation and entropy regularization. By contrast, DDPG-MFTG diverges, relying on a mean-field oracle, which is valid only in the infinite-population limit, and “central players” that map mean-field distributions to deterministic policies without clipping or regularization, making it unstable. Unlike multi-agent DDPG extensions (e.g., MADDPG), which consider other teams’ local policies, DDPG-MFTG conditions only on its own, limiting inter-team awareness. Figure 2(b) shows MF-MAPPO’s scalability, where larger populations reduce noise and variance, aligning with Theorem 3. MF-MAPPO updates every $T_{\text{rollout}}$ steps, unlike DDPG-MFTG which updates per time step, making MF-MAPPO significantly faster and efficient (2h vs. 60h).
@@
We compare MF-MAPPO and DDPG-MFTG by pitting them against each other on a $4\times 4$ grid with full MF information. As shown in Figure 3I, MF-MAPPO consistently outperforms DDPG-MFTG across various initializations, achieving up to $10\times$ higher rewards when attacking and lower/comparable rewards while defending.
@@
Figure 3: I. Average test rewards for Battlefield on a 4x4 grid (100 initializations); II. Example configuration; III. Comparing $\mathrm{d}_{\mathrm{TV}}\big(\cdot\big)$ for D-PC and Benchmark estimator for different $R_{\textrm{com}}$.
@@
Figure 3II shows MF-MAPPO Red agents successfully cover both corridors and deactivate several Blue attackers II(a). Panels II(a) and II(b) highlight that DDPG-MFTG Blue agents do not aggressively pursue the target, illustrating their tendency to passively seek zero-reward outcomes rather than take goal-directed actions, unlike II(c) where MF-MAPPO agents exhibit coordinated maneuvering, forming coalitions to reach the target. While cases I and II utilize complete observability, case III evaluates the proposed D-PC estimator against the estimator in Benjamin and Abate, 2025a (Benchmark) under a gradient-penalized MF-MAPPO policy when Red has full information and Blue estimates Red’s distribution. Both estimators yield comparable total variation errors relative to the full-information case, with D-PC showing advantages under low communication budgets ($R_{\textrm{com}}<20$). One can trivially show that the Benchmark satisfies Proposition 3 with $\epsilon=1-\mathcal{O}(R_{\textrm{com}}/|\mathcal{X}|)$ and assumes uniform estimates for unobserved states, which degrades estimation accuracy under limited communication. It also relies on accurate information from neighbors to ensure validity of estimates. In contrast, D-PC exchanges inexact information but applies state-dependent corrections (projection), preserving privacy and robustness.
@@
In Figure 4, the Red team faces a dilemma in deciding which target to defend, while the Blue team exploits this ambiguity. Due to DDPG-MFTG’s high computational cost and frequent network updates, it is excluded from our analysis. With no other baselines for such large-scale complex MFTGs, we qualitatively assess MF-MAPPO’s performance. Figures 4I–II illustrate how identical policies can generate heterogeneous team behaviors, with Blue adapting target selection and Red reallocating defenses, highlighting the flexibility of the mean-field approximation. Furthermore, D-PC again outperforms the Benchmark under limited communication (III) ($R_{\textrm{com}}<10$) and performs competitively otherwise. Cumulative rewards (IV) show only small deviations, consistent with Proposition 3 and Theorem 4, confirming that agents can rely on local observations with minimal communication.
@@
Figure 4: I. Red is concentrated; 30% Blue are at the bottom, rest are at the top II. Blue is evenly split, Red is concentrated III. Comparing $\mathrm{d}_{\mathrm{TV}}\big(\cdot\big)$ for D-PC and the Benchmark estimator for different communication rounds IV. % error in cumulative rewards: D-PC vs. fully observable ZS-MFTG.
@@ Appendix A Details of D-PC
To satisfy Assumption 4, one may use the well-known Metropolis Matrix [Xiao et al.,, 2006]. We define the smallest non-zero entry of $\alpha$ as
@@
One could alternatively cast this problem as a Partially Observable Markov Decision Process (PO-MDP) [Bernstein et al.,, 2002], where the environment provides arbitrary observations. However, our focus is on policies that explicitly depend on the mean-field distributions of the two teams, for two key reasons (1) Distinct opponent strategies induce different mean-field trajectories, and closed-loop feedback on these distributions enables an appropriate response to the strategies deployed by the opponent [Guan et al., 2024b, ] and (2) Access to team-level distributions facilitates credit assignment [Pignatelli et al.,, 2024] in MARL by allowing agents to reason about which collective distributions are optimal. Thus, having access to MF information—particularly that of the opponent—is desirable. Yet, it is unrealistic to assume that the environment directly provides these distributions, which motivates the design of estimation algorithms. Finally, unlike general PO-MDP formulations that often require maintaining long observation histories, our setting involves population sizes for which such history-based tracking is computationally infeasible in terms of memory.
@@ Appendix B Proof of Theoretical Results
In MF-MAPPO we directly train the finite-population local policies $(\phi_{t},\psi_{t})$. As a result, the gradient of the coordinator policy $\nabla_{\theta}J(\alpha_{\theta})$ only has action samples $u^{N_{1}}_{i,t},i\in[N_{1}]$ from the finite-population local policies $\phi_{t}(\cdot|x^{N_{1}}_{i,t},\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ and not $\pi^{N_{1}}_{t}(\cdot|x^{N_{1}}_{i,t})\sim\alpha(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ itself. Thus, our first step is to construct an approximation of $\pi^{N_{1}}_{t}$ using the obtained finite-population action samples $u^{N_{1}}_{i,t},i\in[N_{1}]$, resulting in an approximate finite-population policy gradient expression given by $\nabla_{\theta}\hat{J}^{N_{1}}(\alpha_{\theta})$. Coupled with existing mean-field approximation results [Cui et al.,, 2023; Guan et al., 2024b, ; Shao et al., 2024a, ] for the infinite-population limit $N_{1},N_{2}\to\infty$ (LLN, etc.) we show convergence. ∎
@@
From Guan et al., 2024b it follows that there exists a one-to-one correspondence between the deterministic infinite-population coordinator policies and the local identical team policies followed by the finite-population Blue and Red agents. We extend this formulation to potentially stochastic infinite-population coordinator policies and the local finite-population policies as follows: a stochastic Blue coordination policy $\alpha\in\mathcal{A}$ induces an identical team policy $\phi\in\Phi$ according to the rule
@@
It is pertinent to note that the action-space $\Pi$ and $\Sigma$ are continuous. Thus, although Assumption 5 seems restrictive, the use of function approximators to train $Q^{\alpha,\beta}$ and $Q^{N,\alpha,\beta}$ like neural networks (endowed with continuous activation functions) in such continuous action-space settings, ensures that the resultant functions are continuous in their inputs Lillicrap et al., [2019].
The following assumption, as done in many works [Cui et al.,, 2024, 2023], states that the gradient of coordinator policy is continuous and uniformly bounded, i.e.,
@@
Furthermore, we have the following result from Guan et al., 2024b :
@@
Assumption 1 allows the $Q^{N,\alpha,\beta},Q^{\alpha,\beta}$ to have compact support. Thus, it suffices to prove pointwise convergence [Cui et al.,, 2023] for each $\tau$, i.e.,
@@
as $N_{1},N_{2}\to\infty$, where the expectation is taken with respect to $\alpha,\beta,\mathbf{u}\sim\pi,\mathbf{v}\sim\sigma$ (dropped for readability). We follow an induction approach similar to the one presented in Cui et al., [2023]. It trivially holds at $\tau=0$. Let us assume that (B) holds at time step $\tau$. For subsequent $\tau+1$,
@@
The convergence of the first term follows from the weak LLN Cui et al., [2023] and a similar argument as used for (B.19). By using the Lipshcitzness of the reward function (Assumption 1) and Corollary 1, the second term in the expression is bounded by $\mathcal{O}\Big(\frac{1}{\sqrt{\underline{N}}}\Big)$, where $\underline{N}=\min(N_{1},N_{2})$ and goes to zero as $(N_{1},N_{2})\to\infty.$
@@
The key idea of the proof is to exploit the fact that the equivalent infinite-population games for $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are the same as $N_{1}/N_{2}=\bar{N}_{1}/\bar{N}_{2}$ [Guan et al., 2024b, ]. We first compute the performance of $\phi^{*}$, the optimal identical team policy derived for the $\mathcal{G}_{1}$ using MF-MAPPO (finite-population training), when applied to the infinite-population coordinator setting. We then compare the performance of the aforementioned $\phi^{*}$ for $\mathcal{G}_{2}$ with the infinite-population coordinator. Then, equating the common equivalent coordinators for $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, the result follows. ∎
@@
For inequality (i), we used the inductive hypothesis; for inequality (ii), we reduced the optimization domain of the Red team to the set of identical team policies; inequality (iv) is a result of the Lipschitz continuity of the value function [Guan et al., 2024b, ]; for equality (v), we use the fact that the mean-fields $(\mu^{\phi^{*}}_{t+1},\nu^{\psi_{t}}_{t+1})$ are induced deterministically from the distributions at time $t$; inequality (vi) holds as a consequence of Corollary 1; (vii) converts the optimization domain from the policy space $\psi_{t}\in\Psi_{t}$ to the corresponding reachable set $\mathds{R}_{\nu,t}(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})\triangleq\{\nu_{t+1}|\exists\psi_{t}\in\Psi_{t}~\mathrm{s.t.}~\!\nu_{t+1}=\nu_{t}G_{t}(\mu_{t},\nu_{t},\psi_{t})\}$ following [Guan et al., 2024b, ]; (vii) follows from the definition of $J_{\infty,t}^{\phi^{*},\tilde{\psi}^{*}_{\infty}}(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$. Thus, $J^{N,\phi^{*},\psi^{N_{2}*}}(\mathbf{x}^{N_{1}},\mathbf{y}^{N_{2}})=J^{N,\phi^{*},\psi^{N_{2}*}}_{0}(\mathbf{x}^{N_{1}},\mathbf{y}^{N_{2}})$ and (B.22) follows.
@@
For the identical team policy $\phi^{*}\in\Phi$, denote $\mu^{\phi^{*}}_{t+1})$ from (B). Furthermore, from Theorem 1 in Guan et al., 2024b there exists a $\nu^{\psi^{\bar{N}_{2}}_{t}}_{\mathrm{apprx},t+1}$ for the Red team policy $\psi_{t}^{N_{2}}$ such that

### v3

@@ 4 MF-MAPPO for Partially Observable MFTGs
Similarly, we can define Lipschitz continuous policies for the Red team with constant $L_{\psi}.$ This idea of penalizing the gradient of the policies was introduced in robotics to promote smooth and stable policies in order to aid sim-to-real transfer [Chen et al., 2024; Shin et al., 2025].
@@
In lieu of Proposition 3, we propose a communication network-based decentralized estimation filter, namely, Dynamic-Projected Consensus (D-PC). It is an extension of the control-theoretic constrained consensus problem [Nedić and Liu, 2016] and addresses the shortcomings of the estimation algorithm proposed in Benjamin and Abate [2025b], namely, estimation under limited communication rounds, and ensuring valid estimates in the presence of errors. Following the connected graph topology used in MFGs [Benjamin and Abate, 2025b], we define a state-based visibility graph $\mathcal{G}_{t}^{\mathrm{viz}}$ and team-based communication graphs $\mathcal{G}_{{\rm Blue},t}^{\mathrm{com}}$ and $\mathcal{G}_{{\rm Red},t}^{\mathrm{com}}$. We also define the projection operator $\Omega_{\mathcal{R}(x)}[\eta]\triangleq\arg\min_{\omega\in\mathcal{R}(x)}\|\eta-\omega\|_{2}$ for $\eta,\omega\in\mathbb{R}^{|\mathcal{Y}|}$ where $\mathcal{R}(x)$ is a closed and convex set. We assume that all agents in the same state receive the same information, so naturally they have the same estimate, i.e., all Blue agents at state $x\in\mathcal{X}$ at time $t$ have the same estimate $\hat{\nu}^{N_{2}}_{x,t}$ of the Red team MF. We consider two time scales: $t$ for system dynamics (2) and $\tau$ for communication rounds. At time $t$, and $\tau=0$, each Blue agent at state $x\in\mathcal{X}$ holds a belief $\hat{\nu}^{\tau=0}_{x,t}$ consistent with $\mathcal{G}_{t}^{\mathrm{viz}}$. At communication round $\tau$, agents share estimates $\hat{\nu}^{\tau-1}_{x,t}$ with neighbors defined by $\mathcal{G}_{{\rm Blue},t}^{\mathrm{com}}$ and perform for $R_{\textrm{com}}$ communication rounds: a) weighted-average consensus; and b) a projection onto a closed and convex constraint set $\mathcal{R}(x)$. The set $\mathcal{R}(x)$ combines the Red team’s MF components known with certainty by the Blue agents at state xx—i.e., the observable states given by $\mathcal{G}_{t}^{\mathrm{viz}}$—with those that must be estimated. $\mathcal{R}(x)$ guarantees that operations such as information aggregation, averaging, or distributed communication do not alter the parts of the distribution that are known with certainty. See Appendix A for detailed definitions.
@@ 5 Numerical Experiments
In this section, we evaluate MF-MAPPO across large-population scenarios using our custom-made benchmark simulation platform, MFEnv. Built as an extension of Gymnasium [Towers et al., 2024], MFEnv is developed specifically to facilitate research in MFTGs, supporting both finite-agent simulations and oracle-based infinite-population models. Unlike existing toolkits, MFEnv includes aggregate reward metrics, policy-versus-policy evaluation, and flexible APIs for custom mean-field environments that adhere to MF dynamics, rewards and information structures. We showcase MF-MAPPO’s efficacy on three representative environments (1) a constrained-action variant of the classical rock–paper–scissors game [Raghavan, 1994], enabling validation against analytically computed equilibria, (2) a complex battlefield setting where Blue and Red teams engage in attack–defense tasks with higher-dimensional state and action spaces, requiring sophisticated team-level coordination (3) a ZS-MFTG between a virus and a population modeling a SIS-framework. Additional results and environments (not limited to ZS-MFTGs) in Appendix E.
To the best of our knowledge, DDPG-MFTG [Shao et al., 2024a] is the only existing algorithm explicitly designed for mixed collaborative-competitive mean-field team problems. Accordingly, we mainly benchmark against DDPG-MFTG as well as several “house-made” mean-field adaptations of established MARL methods: namely, independent PPO (MF-IPPO) and two variants of MAPPO—parameter sharing (MAPPO-PS) and centralized critic (MAPPO-CC) [Yu et al., 2022]. Centralized Q-learning methods such as QMix [Rashid et al., 2020], MADDPG [Lowe et al., 2017], and FACMAC [Peng et al., 2021] are not suitable baselines in our setting because their centralized action-value functions depend on the joint action and full system state, making them intractable in large-population regimes. Moreover, these approaches yield deterministic local policies, which tend to perform poorly since mixed strategies produce the heterogeneous team behaviors essential for mission success, as shown in Figure 4, making such comparisons inherently unfair.
@@
Constrained Rock-Paper-Scissors (cRPS). The state space of each individual agent is $\mathcal{S}=\{\texttt{R,P,S}\}$, representing rock, paper, and scissors, respectively. We consider a non-trivial restriction of the action space to $\mathcal{A}=\{\texttt{CW},\texttt{Stay}\}$ allowing agents to either move clockwise ( $\texttt{R}\rightarrow\texttt{P}$, $\texttt{P}\rightarrow\texttt{S}$, $\texttt{S}\rightarrow\texttt{R}$) or remain idle, respectively. We assume deterministic transitions, where each action leads to a unique next state deterministically. At each time step $t$, the Blue (Red) team receives a team reward $r_{t}(\mu_{t},\nu_{t})={\mu}_{t}^{\mbox{\sf T}}A{\nu}_{t}$ (${\mu}_{t}^{\mbox{\sf T}}A{\nu}_{t}$) where $A$ is the standard RPS payoff matrix. Table 1 reports performance comparisons. MF-MAPPO achieves rewards closest to the analytical Nash value (1/3)(-1/3). Its minimally-informed critic preserves local-state privacy while substantially reducing critic dimensionality, resulting in a faster and more efficient algorithm under identical operating conditions. Figure 2(a) compares trajectories from MF-MAPPO and DDPG-MFTG. We see that MF-MAPPO successfully reaches the equilibrium distribution. By contrast, DDPG-MFTG diverges, relying on a mean-field oracle, which is valid only in the infinite-population limit, and “central players” that map mean-field distributions to deterministic policies without clipping or regularization, making it unstable. Unlike multi-agent DDPG extensions (e.g., MADDPG), which consider other teams’ local policies, DDPG-MFTG conditions only on its own, limiting inter-team awareness. Figure 2(b) shows MF-MAPPO’s scalability, where larger populations reduce noise and variance, aligning with Theorem 3.
Table 1: Performance comparison for cRPS
Algorithm Critic Input Average Reward NE Attained? Training Time
MF-IPPO $\{(x^{N_{1}}_{i,t},\mu_{t})\}_{i=1}^{N_{1}}$ -0.340 2:45:11
MAPPO-PS $\{(x^{N_{1}}_{i,t},\mu_{t},\nu_{t})\}_{i=1}^{N_{1}}$ -0.313 2:47:54
MAPPO-$C$ $(\mathbf{x}^{N_{1}}_{t},\mathbf{y}^{N_{2}}_{t},\mu_{t},\nu_{t})$ -0.342 3:20:46
DDPG-MFTG $(\phi_{t},\mu_{t},\nu_{t})$ 3.774 60:49:35
MF-MAPPO $(\mu_{t},\nu_{t})$ -0.331 2:17:15
@@
We compare MF-MAPPO with the baselines by evaluating them head-to-head on a $4\times 4$ grid with full MF information (Table 2). Teams trained with MF-MAPPO demonstrate consistently superior performance. Fixing the Red team policy (row) shows that the MF-MAPPO Blue team achieves the highest rewards, and fixing the Blue team policy (column) shows that the MF-MAPPO Red team performs the strongest as the minimizing team. These results highlight the benefits of using a minimally informed critic together with a shared-team actor: the resulting networks are smaller, more scalable, preserve private information, and alleviate credit-assignment challenges.
@@
Figure 3: I. Training curve for Battlefield on a 4x4 grid (Blue team); II. Example configuration; III. Comparing $\mathrm{d}_{\mathrm{TV}}\big(\cdot\big)$ for D-PC and Benchmark estimator for different $R_{\textrm{com}}$.
@@
Figure 3II shows MF-MAPPO Red agents successfully cover both corridors and deactivate several Blue attackers II(a). Panels II(a) and II(b) highlight that DDPG-MFTG Blue agents do not aggressively pursue the target, illustrating their tendency to passively seek zero-reward outcomes (also reflected in Panel I) rather than take goal-directed actions, unlike II(c) where MF-MAPPO agents exhibit coordinated maneuvering, forming coalitions to reach the target. While cases I and II utilize complete observability, case III evaluates the proposed D-PC estimator against the estimator in Benjamin and Abate [2025a] (Benchmark) under a gradient-penalized MF-MAPPO policy when Red has full information and Blue estimates Red’s distribution. Both estimators yield comparable total variation errors relative to the full-information case, with D-PC showing advantages under low communication budgets ($R_{\textrm{com}}<20$). One can trivially show that the Benchmark satisfies Proposition 3 with $\epsilon=1-\mathcal{O}(R_{\textrm{com}}/|\mathcal{X}|)$ and assumes uniform estimates for unobserved states, which degrades estimation accuracy under limited communication. It also relies on accurate information from neighbors to ensure validity of estimates. In contrast, D-PC exchanges inexact information but applies state-dependent corrections (projection), preserving privacy and robustness.
Table 2: Performance evaluation for Battlefield
MF-IPPO MAPPO-PS MAPPO-$C$ DDPG-MFTG MF-MAPPO
MF-IPPO 61.54 73.85 70.60 0.00 86.26
MAPPO-PS 55.82 72.78 71.64 0.00 86.22
MAPPO-$C$ 51.60 68.81 65.58 0.00 82.58
DDPG-MFTG 65.26 75.14 77.21 0.00 85.95
MF-MAPPO 36.10 54.43 62.00 0.00 76.26
@@
In Figure 4, the Red team faces a dilemma in deciding which target to defend, while the Blue team exploits this ambiguity. Due to DDPG-MFTG’s high computational cost and frequent network updates, it is excluded from our analysis. With no other baselines for such large-scale complex MFTGs, we qualitatively assess MF-MAPPO’s performance. Figures 4I–II illustrate how identical policies can generate heterogeneous team behaviors, with Blue adapting target selection and Red reallocating defenses, highlighting the flexibility of the mean-field approximation. Furthermore, D-PC again outperforms the Benchmark under limited communication (III) ($R_{\textrm{com}}<10$) and performs competitively otherwise. Cumulative rewards (IV) show only small deviations, consistent with Proposition 3 and Theorem 4, confirming that agents can rely on local observations with minimal communication, and that performance improves as communication bandwidth increases. We further empirically evaluate the effect of communication quality on opponent-estimate performance (V). Fixing $R_{\textrm{com}}=7$, injecting zero-mean Gaussian noise with increasing variance produces larger deviations in cumulative rewards; however, these remain within $5\%$ tolerance, due to the projection step after the average-consensus update, ensuring robust performance.
@@
Figure 4: I. Red is concentrated; 30% Blue are at the bottom, rest are at the top II. Blue is evenly split, Red is concentrated III. Comparing $\mathrm{d}_{\mathrm{TV}}\big(\cdot\big)$ for D-PC and the Benchmark estimator for different $R_{\textrm{com}}$ IV. % error in cumulative rewards under varying communication bandwidths V. % error in cumulative rewards under noisy communication $R_{\textrm{com}}=7.$
Epidemiology. In this scenario, the Red agents emulate a virus and can infect the Blue population; infected Blue individuals can further transmit the disease. Blue agents aim to maximize the size of their healthy population by avoiding infected peers. Because the virus cannot be eliminated, the infected Blue agents must visit a hospital state, where healing occurs with some probability. Applying MF-MAPPO to this ZS-MFTG yields intuitive policies: decongestion-like behavior in which agents spread out to reduce contact (Figure 5). Meanwhile, the Red team actively tracks Blue agents to infect them and promptly positions itself around the hospital, while infected Blue agents strategically navigate toward the hospital for treatment.
Figure 5: Example configurations of SIS-Epidemiology ZS-MFTG; green cell is the hospital.
@@ Appendix A Details of D-PC
To satisfy Assumption 4, one may use the well-known Metropolis Matrix [Xiao et al., 2006]. We define the smallest non-zero entry of $\alpha$ as
@@
One could alternatively cast this problem as a Partially Observable Markov Decision Process (PO-MDP) [Bernstein et al., 2002], where the environment provides arbitrary observations. However, our focus is on policies that explicitly depend on the mean-field distributions of the two teams, for two key reasons (1) Distinct opponent strategies induce different mean-field trajectories, and closed-loop feedback on these distributions enables an appropriate response to the strategies deployed by the opponent [Guan et al., 2024b] and (2) Access to team-level distributions facilitates credit assignment [Pignatelli et al., 2024] in MARL by allowing agents to reason about which collective distributions are optimal. Thus, having access to MF information—particularly that of the opponent—is desirable. Yet, it is unrealistic to assume that the environment directly provides these distributions, which motivates the design of estimation algorithms. Finally, unlike general PO-MDP formulations that often require maintaining long observation histories, our setting involves population sizes for which such history-based tracking is computationally infeasible in terms of memory.
@@ Appendix B Proof of Theoretical Results
In MF-MAPPO we directly train the finite-population local policies $(\phi_{t},\psi_{t})$. As a result, the gradient of the coordinator policy $\nabla_{\theta}J(\alpha_{\theta})$ only has action samples $u^{N_{1}}_{i,t},i\in[N_{1}]$ from the finite-population local policies $\phi_{t}(\cdot|x^{N_{1}}_{i,t},\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ and not $\pi^{N_{1}}_{t}(\cdot|x^{N_{1}}_{i,t})\sim\alpha(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ itself. Thus, our first step is to construct an approximation of $\pi^{N_{1}}_{t}$ using the obtained finite-population action samples $u^{N_{1}}_{i,t},i\in[N_{1}]$, resulting in an approximate finite-population policy gradient expression given by $\nabla_{\theta}\hat{J}^{N_{1}}(\alpha_{\theta})$. Coupled with existing mean-field approximation results [Cui et al., 2023; Guan et al., 2024b; Shao et al., 2024a] for the infinite-population limit $N_{1},N_{2}\to\infty$ (LLN, etc.) we show convergence. ∎
@@
From Guan et al. [2024b] it follows that there exists a one-to-one correspondence between the deterministic infinite-population coordinator policies and the local identical team policies followed by the finite-population Blue and Red agents. We extend this formulation to potentially stochastic infinite-population coordinator policies and the local finite-population policies as follows: a stochastic Blue coordination policy $\alpha\in\mathcal{A}$ induces an identical team policy $\phi\in\Phi$ according to the rule
@@
It is pertinent to note that the action-space $\Pi$ and $\Sigma$ are continuous. Thus, although Assumption 5 seems restrictive, the use of function approximators to train $Q^{\alpha,\beta}$ and $Q^{N,\alpha,\beta}$ like neural networks (endowed with continuous activation functions) in such continuous action-space settings, ensures that the resultant functions are continuous in their inputs Lillicrap et al. [2019].
The following assumption, as done in many works [Cui et al., 2024, 2023], states that the gradient of coordinator policy is continuous and uniformly bounded, i.e.,
@@
Furthermore, we have the following result from Guan et al. [2024b]:
@@
Assumption 1 allows the $Q^{N,\alpha,\beta},Q^{\alpha,\beta}$ to have compact support. Thus, it suffices to prove pointwise convergence [Cui et al., 2023] for each $\tau$, i.e.,
@@
as $N_{1},N_{2}\to\infty$, where the expectation is taken with respect to $\alpha,\beta,\mathbf{u}\sim\pi,\mathbf{v}\sim\sigma$ (dropped for readability). We follow an induction approach similar to the one presented in Cui et al. [2023]. It trivially holds at $\tau=0$. Let us assume that (B) holds at time step $\tau$. For subsequent $\tau+1$,
@@
The convergence of the first term follows from the weak LLN Cui et al. [2023] and a similar argument as used for (B.19). By using the Lipshcitzness of the reward function (Assumption 1) and Corollary 1, the second term in the expression is bounded by $\mathcal{O}\Big(\frac{1}{\sqrt{\underline{N}}}\Big)$, where $\underline{N}=\min(N_{1},N_{2})$ and goes to zero as $(N_{1},N_{2})\to\infty.$
@@
The key idea of the proof is to exploit the fact that the equivalent infinite-population games for $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are the same as $N_{1}/N_{2}=\bar{N}_{1}/\bar{N}_{2}$ [Guan et al., 2024b]. We first compute the performance of $\phi^{*}$, the optimal identical team policy derived for the $\mathcal{G}_{1}$ using MF-MAPPO (finite-population training), when applied to the infinite-population coordinator setting. We then compare the performance of the aforementioned $\phi^{*}$ for $\mathcal{G}_{2}$ with the infinite-population coordinator. Then, equating the common equivalent coordinators for $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, the result follows. ∎
@@
For inequality (i), we used the inductive hypothesis; for inequality (ii), we reduced the optimization domain of the Red team to the set of identical team policies; inequality (iv) is a result of the Lipschitz continuity of the value function [Guan et al., 2024b]; for equality (v), we use the fact that the mean-fields $(\mu^{\phi^{*}}_{t+1},\nu^{\psi_{t}}_{t+1})$ are induced deterministically from the distributions at time $t$; inequality (vi) holds as a consequence of Corollary 1; (vii) converts the optimization domain from the policy space $\psi_{t}\in\Psi_{t}$ to the corresponding reachable set $\mathds{R}_{\nu,t}(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})\triangleq\{\nu_{t+1}|\exists\psi_{t}\in\Psi_{t}~\mathrm{s.t.}~\!\nu_{t+1}=\nu_{t}G_{t}(\mu_{t},\nu_{t},\psi_{t})\}$ following [Guan et al., 2024b]; (vii) follows from the definition of $J_{\infty,t}^{\phi^{*},\tilde{\psi}^{*}_{\infty}}(\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$. Thus, $J^{N,\phi^{*},\psi^{N_{2}*}}(\mathbf{x}^{N_{1}},\mathbf{y}^{N_{2}})=J^{N,\phi^{*},\psi^{N_{2}*}}_{0}(\mathbf{x}^{N_{1}},\mathbf{y}^{N_{2}})$ and (B.22) follows.
@@
For the identical team policy $\phi^{*}\in\Phi$, denote $\mu^{\phi^{*}}_{t+1})$ from (B). Furthermore, from Theorem 1 in Guan et al. [2024b] there exists a $\nu^{\psi^{\bar{N}_{2}}_{t}}_{\mathrm{apprx},t+1}$ for the Red team policy $\psi_{t}^{N_{2}}$ such that